What is the empirical formula of a compound composed of 32.5 g potassium (K) and 6.65 g oxygen (O)?
1
Expert's answer
2012-11-14T08:34:44-0500
n (K) = 32.5/39 = 0.833 moln(O) = 6.65/16 = 0.415 mol
n(K) : n(O) = 2:1

Empirical formula: K2O